OpenStudy (anonymous):

In a particular region of a national park, there are currently 330 deer, and the population is increasing at an annual rate of 11%. a. Write an exponential function to model the deer population. b. Explain what each value in the model represents. c. Predict the number of deer that will be in the region after five years. Show your work.

OpenStudy (anonymous):

As the population of deer are increasing at an annual rate of 11%, the total population at the end of each year would be 100% + 11%=111%=1.11 as a decimal. So you multiply the population at the end of each year by 1.11 to get 111%.

OpenStudy (anonymous):

so its about 556 deer after 5 years?

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

You're welcome. Did you get the exponential model?

OpenStudy (anonymous):

I think so: y=330*1.11^n y=330*1.11^5 y=330*1.685 y=556.05

OpenStudy (anonymous):

Good work! Oh, and you shouldn't need to round 1.11^5 if you put the 330*1.11^5 straight into your calculator - assuming it's a scientific one.

OpenStudy (anonymous):

thank you sooooo much!!!

OpenStudy (anonymous):

No prob!